Evaluate the derivative of the following function. f(w) = sin [cos - (7w)] f(w)=
Answer #1

f(w) = sin sis[cos-(760)] Formulla sin (los) Nina2 2 fiws=Ni-(70) ² Now, !!20) = d (1-49W w?) un dw 6-1 id (1-49 6?) = (lae9